OpenStudy (anonymous):

How many solutions does the system of equation have? 3x=-12y+15 and x+4y=5

OpenStudy (anonymous):

Ok so here are both of the equations graphed

OpenStudy (anonymous):

That would make the answer 2?

OpenStudy (anonymous):

(By the way, you can solve these by Substitution,Elimination, or Graphing, i'm solving by graph because it's the easiest right now)

OpenStudy (anonymous):

It touches both y and x 1 time in the graph

OpenStudy (anonymous):

The lines are on top of each other. You know that parallel lines have no solutions, how many do you think this has?

OpenStudy (anonymous):

Ohhhhh so... the solution is how many times the 2 LINES touch each other; not the x or y axis?

OpenStudy (anonymous):

When you do Substitution or Elimination, you're solving for the x and y variables. Those x and y variables are coordinates on where the 2 lines touch each other (x,y)

OpenStudy (anonymous):

So it is infinitely because they are virtually the same line

OpenStudy (anonymous):

Yes! =)

OpenStudy (anonymous):

x-4y=12 and 5x-20y=60

OpenStudy (anonymous):

Infinitely again

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

You can tell it's infinite because if you do elimination, multiply the first equation by 5 and you see that they resemble each other and when subtracted will equal 0.

OpenStudy (anonymous):

Ok I see that now. Thx!! y-6x=-3 and 4y-24x=-16 (last one)

OpenStudy (anonymous):

would you multiply every digit by 24? How do you find the number of what to multiply the equations by?

OpenStudy (anonymous):

No, you would just multiply the first equation by 4 so that the -6x turns into -24x, then subtract them to cancel out the x's and solve for y.

OpenStudy (anonymous):

Oh! I see now!

OpenStudy (triciaal):

the objective is to make the coefficient of one of the variable equal and opposite

OpenStudy (anonymous):

what is the coefficient? What's the definition of that?

OpenStudy (anonymous):

2x, the coefficient is 2 because it's infront of the x

OpenStudy (anonymous):

and that rule applies ALWAYS?

OpenStudy (anonymous):

And here is the graph to the last question you asked so you can check your ans

OpenStudy (anonymous):

the 4 is the slope of the line? Correct?

OpenStudy (anonymous):

Yes! I got it right! Thx!

OpenStudy (anonymous):

Np, goodluck on any other =)

OpenStudy (anonymous):

I understand that topic now

OpenStudy (triciaal):

one of the 4 ways to solve simultaneous equations is the elimination/addition method. you eliminate one of the variable to get a linear equation. you have one unknown. solve for that variable then substitute the value in either of the original expressions for the other variable and get the other value.

OpenStudy (anonymous):

Ok thnx!

OpenStudy (anonymous):

I will be posting just a few more (different) questions

OpenStudy (triciaal):

it is always a very good practice to check the solution because it is very very common to make mistakes (arithmetic error)